OpenStudy (anonymous):

simplify 4[3-(1-6)^3]

hero (hero):

4[3 - (-5)^3] 4[3 +125] 4[128] = 512

OpenStudy (anonymous):

please go tell your math teacher that the word is not "simplify"

hero (hero):

Evaluate

OpenStudy (anonymous):

compute maybe find the number what is ___? or as hero said evaluate god not wonder people get confused

OpenStudy (anonymous):

When you have a variable then "simplify" would be a appropriate.

OpenStudy (anonymous):

even Find

OpenStudy (anonymous):

@brandon like hell

OpenStudy (anonymous):

Haha

OpenStudy (anonymous):

what does "simplify" \[(x-2)(x+3)\] mean? nothing

OpenStudy (anonymous):

but math teachers write it when they mean "multiply"

OpenStudy (anonymous):

or what does "simplify" \[x^2-2x+3x-10\] mean? it means "combine like terms"

OpenStudy (anonymous):

or "simplify"\[\frac{x^2-1}{x-1}\] it means factor and reduce to lowest terms

OpenStudy (anonymous):

\[\frac{x^{2}+12-36}{2x+8}*\frac{x^{2}+56-90}{6x+12}\] I would call reducing expressions like these "simplifying"

OpenStudy (anonymous):

they use it to mean "write what i want"

OpenStudy (anonymous):

ok but i would stop with the word "reduce" and forget about the word "simplify"

OpenStudy (anonymous):

I see what you mean though it is used in appropriate places. If there's no variable then there is no need for it.

OpenStudy (anonymous):

really teachers should say what they mean. there is nothing more confusing than using one word to mean may different things. especially in math

OpenStudy (anonymous):

Hey satellite 4 is the simplified form of 2+2 :O)

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

i cringe when i see "simplify" \[\frac{8}{12}\]

OpenStudy (anonymous):

suppose my job is to add \[\frac{1}{4}+\frac{2}{3}\]? much "simpler' if i have \[\frac{3}{12}+\frac{8}{12}\]